# On the Salient Limitations of the Methods of Assembly Theory and their Classification of Molecular Biosignatures

### 10 Jan 2024 v6

@@ A.2.2 Definitions and notation
Let $\gamma=z\dots y$ denote an arbitrary path from $z\in B_{\mathcal{S}}$ to some $y\in V\left(\mathcal{S}\right)$ in $\mathcal{S}$, where $B_{\mathcal{S}}$ is the basis (i.e., the finite set of basic building blocks) of $\mathcal{S}$ and $V\left(\mathcal{S}\right)$ is the set of vertices of $\mathcal{S}$. Let $\gamma_{x}$ denote a path from some $z\in B_{\mathcal{S}}$ to the object $x\in V\left(\mathcal{S}\right)$.
Let ${\Gamma^{*}}_{x}$ denote a minimum rooted assembly subspace of $\Gamma$ from which the assembly index $c_{\Gamma}\left(x\right)$ calculates the augmented cardinality and that its longest rooted path $\gamma_{x}$ ends in the arbitrary object $x\in V\left(\Gamma\right)$ as in [Marshall2019, Definition 19].
@@
As assumed to the above, and in accordance with the claims in [cronin, marshall_murray_cronin_2017, Marshall2019], notice that the computability and feasibility of assembly theory’s methods directly implies the existence of such a submachine. Given any set of (physical, chemical, and/or biological) rules for assembling molecules that assembly theory arbitrarily chooses to be based on and that is consistent with our current knowledge of physics, one has it that for any particular pathway assembly resulting in a molecule, there is a correspondent program running on a submachine $\mathbf{U}_{\Gamma}$ for which one can apply the methods described in [cronin, marshall_murray_cronin_2017, Marshall2019] to calculate the pathway probability and so on. This existence is trivially guaranteed to hold as long as the calculation of the assembly index is computable, which is already an assumed condition in [cronin, marshall_murray_cronin_2017, Marshall2019]; moreover, any upper bound for the computational resources necessary to compute the assembly index straightforwardly implies the existence of an upper bound for the computational resources necessary for a program to run on $\mathbf{U}_{\Gamma}$ (i.e., submachine $\mathbf{U}_{\Gamma}$ is a type of resource-bounded Turing machine).
@@
For the sake of simplifying notation, $\mathbf{U}_{\Gamma}$ denotes the Turing submachine $\mathbf{U}/{f}_{\Gamma}$. In this case, the function $f_{\Gamma}$ is the partial function that returns what $\mathbf{U}$ can compute with some $x\in\mathbf{L}_{\Gamma}$ as input, where $\mathbf{L}_{\Gamma}\subseteq\mathbf{L}$ is a (non-universal) programming language such that every allowed—i.e., deemed physically possible by assembly theory’s chosen method, in case the objects are molecules—generative process of an assembly space is bijectively computed (or emulated) by a corresponding $\mathbf{U}\left(x\right)$. In other words, for every allowed generative process that can assemble objects into building another object, there is a program $x\in\mathbf{L}_{\Gamma}$ that computes (or emulates) this process. Conversely, for every $x\in\mathbf{L}_{\Gamma}$, one also has it that there is a corresponding generative process which is computed (or emulated) by program $x\in\mathbf{L}_{\Gamma}$. Notice that in the forthcoming Theorem A.4, we will show that if a certain generative process is equivalent (in terms of complexity and computation capability) to a particular type of algorithm running on this resource-bounded submachine, then the deceiving phenomenon takes places. Although our theoretical results prove the existence of false positives in the case this condition is met, our empirical results shown in the main article not only corroborate this prediction but in fact suggest this theoretical finding (or some weaker version of it) may be happening in much lower computational classes. This occurs because the assembly index either underperforms or does not outdistance other measures of a statistical nature, where distortions in complexity estimations are already well known to occur—which in particular is one of the reasons complexity science moved on from simplistic statistical measures and remains an evolving field of research.
@@ A.2.3 Theoretical results
The main idea to achieve the following theoretical results is to construct a randomly generated program that receives a formal theory (which contains all the computable procedures and statistical criteria in assembly theory) as input. Then, it searches for a molecule (or object in an assembly space) with MA sufficiently high so as to make the pathway probability of spontaneous formation be sufficiently lower than the very own deceiving program’s algorithmic probability, so that the divergence between these two probability distributions become statistically significant according to the chosen statistical method and significance level.
@@
Let $\mathcal{S}$ be infinite computably enumerable. Let $\mathbf{F}$ be an arbitrary formal theory that contains assembly theory, including all the decidable procedures of the chosen method for calculating (or approximating MA) of an object for a nested subspace of $\mathcal{S}$, and the program that decides whether or not the criteria for building the assembly spaces are met. Let $k\in\mathbb{N}$ be an arbitrarily large natural number. Then, there are a program $\mathrm{p}_{y}$, $\Gamma\in\mathcal{S}$ and $y\in V\left(\Gamma\right)$ such that
@@
holds. In order to demonstrate that $p$ always halts, just note that $\mathcal{S}$ is infinite computably enumerable. Also, for any value of $c_{\Gamma^{\prime}}\left(z\right)$ for some $z\in V\left(\Gamma^{\prime}\right)\subset V\left(\mathcal{S}\right)$, there is only a finite number of possible paths starting on any object in $B_{\mathcal{S}}$ and ending on $z$ in $c_{\Gamma^{\prime}}\left(z\right)$ steps, where $B_{\mathcal{S}}$ is the basis (i.e., the finite set of basic building blocks [cronin, Definition 12]) of $\mathcal{S}$. This implies that there is an infinite number of distinct values of $c_{\Gamma^{\prime}}\left(z\right)$, whenever $z\in V\left(\Gamma^{\prime}\right)$ and $\Gamma^{\prime}\subset\mathcal{S}$ is finite. Now, let $\mathrm{p}_{y}=\left<k,\mathbf{F},p\right>$. Finally, from Equation 4 and basic properties in AIT, we have it that

### 10 Apr 2024 v8

@@ A.2.2 Definitions and notation
Let $\gamma=z\dots y$ denote an arbitrary path from $z\in B_{\mathcal{S}}$ to some $y\in V\left(\mathcal{S}\right)$ in $\mathcal{S}$, where $B_{\mathcal{S}}$ is the basis (i.e., the finite set of basic building blocks) of $\mathcal{S}$ and $V\left(\mathcal{S}\right)$ is the set of vertices of $\mathcal{S}$. Let $\gamma_{x}$ denote a rooted path from some $z\in B_{\mathcal{S}}$ to the object $x\in V\left(\mathcal{S}\right)$.
Let ${\Gamma^{*}}_{x}$ denote a minimum rooted assembly subspace of $\Gamma$ from which the assembly index $c_{\Gamma}\left(x\right)$ calculates the augmented cardinality and that its longest rooted paths $\gamma_{x}$ ends in the arbitrary object $x\in V\left(\Gamma\right)$ as in [Marshall2019, Definition 19].
@@
As assumed to the above, and in accordance with the claims in [cronin, marshall_murray_cronin_2017, Marshall2019], notice that the computability and feasibility of assembly theory’s methods directly implies the existence of such a submachine. Given any set of (physical, chemical, and/or biological) rules for assembling molecules that assembly theory arbitrarily chooses to be based on and that is consistent with our current knowledge of physics, one has it that for any particular pathway assembly resulting in a molecule, there is a correspondent program running on a submachine $\mathbf{U}_{\Gamma}$ for which one can apply the methods described in [cronin, marshall_murray_cronin_2017, Marshall2019] to calculate the pathway probability and so on. This existence is trivially guaranteed to hold as long as the calculation of the assembly index is computable given an assembly space, which is already an assumed condition in [cronin, marshall_murray_cronin_2017, Marshall2019]; moreover, any upper bound for the computational resources necessary to compute the assembly index straightforwardly implies the existence of an upper bound for the computational resources necessary for a program to run on $\mathbf{U}_{\Gamma}$ (i.e., submachine $\mathbf{U}_{\Gamma}$ is a type of resource-bounded Turing machine).
@@
For the sake of simplifying notation, $\mathbf{U}_{\Gamma}$ denotes the Turing submachine $\mathbf{U}/{f}_{\Gamma}$. In this case, the function $f_{\Gamma}$ is the partial function that returns what $\mathbf{U}$ can compute with some $x\in\mathbf{L}_{\Gamma}$ as input, where $\mathbf{L}_{\Gamma}\subseteq\mathbf{L}$ is a (non-universal) programming language such that every allowed generative process—i.e., every generative process that is deemed physically possible by assembly theory’s chosen method, in case the objects are molecules—of an assembly space is bijectively computed (or emulated) by a corresponding $\mathbf{U}\left(x\right)$. In other words, for every allowed generative process that can assemble objects into building another object, there is a program $x\in\mathbf{L}_{\Gamma}$ that computes (or emulates) this process. Conversely, for every $x\in\mathbf{L}_{\Gamma}$, one also has it that there is a corresponding generative process allowed by assembly theory, process which is computed (or emulated) by program $x\in\mathbf{L}_{\Gamma}$. Notice that in the forthcoming Theorem A.4, we will show that if a certain generative process is equivalent (in terms of complexity and computation capability) to a particular type of algorithm running on this resource-bounded submachine, then the deceiving phenomenon takes places. Although our theoretical results prove the existence of false positives in the case this condition is met, our empirical results shown in the main article not only corroborate this prediction but in fact suggest this theoretical finding (or some weaker version of it) may be happening in much lower computational classes. This occurs because the assembly index either underperforms or does not outdistance other measures of a statistical nature, where distortions in complexity estimations are already well known to occur—which in particular is one of the reasons complexity science moved on from simplistic statistical measures and remains an evolving field of research.
@@ A.2.3 Theoretical results
The main idea to achieve the following theoretical results is to construct a randomly generated program that receives a formal theory (which contains all the computable procedures and statistical criteria in assembly theory) as input. Then, it searches for a molecule (or object in an assembly space) with MA sufficiently high so as to make the pathway probability of spontaneous formation be sufficiently lower than the very own deceiving program’s algorithmic probability, so that the divergence between these two probability distributions become statistically significant according to the arbitrarily chosen statistical method and significance level.
@@
Let $\mathcal{S}$ be infinite computably enumerable. Let $\mathbf{F}$ be an arbitrary formal theory that contains assembly theory, including all the decidable procedures of the chosen method for calculating the assembly index (or approximating MA) of an object for a nested subspace of $\mathcal{S}$, and the program that decides whether or not the criteria for building the assembly spaces are met. Let $k\in\mathbb{N}$ be an arbitrarily large natural number. Then, there are a program $\mathrm{p}_{y}$, $\Gamma\subset\mathcal{S}$ and $y\in V\left(\Gamma\right)$ such that
@@
holds. In order to demonstrate that $p$ always halts, just note that $\mathcal{S}$ is infinite computably enumerable. Also, for any value of $c_{\Gamma^{\prime}}\left(z\right)$ for some $z\in V\left(\Gamma^{\prime}\right)\subset V\left(\mathcal{S}\right)$, there is only a finite number of minimum rooted assembly subspaces (starting on any object in $B_{\mathcal{S}}$ and ending on $z$) whose augmented cardinality is $c_{\Gamma^{\prime}}\left(z\right)$, where $B_{\mathcal{S}}$ is the basis (i.e., the finite set of basic building blocks [cronin]) of $\mathcal{S}$. This implies that there is an infinite number of distinct values of $c_{\Gamma^{\prime}}\left(z\right)$. Now, let $\mathrm{p}_{y}=\left<k,\mathbf{F},p\right>$. Finally, from Equation 4 and basic properties in AIT, we have it that